OpenStudy (colcaps):

What is the angular displacement of point A for 15/12 revolutions in radians?

OpenStudy (anonymous):

15/12 revolution = 15/12-1 revolution = 1/4 revolution 1 revolution = 2π radians

OpenStudy (colcaps):

ok thank you